# Post-detection inference for sequential changepoint localization

### v2

@@ D.3.1 Known pre-change and composite post-change (Setting II)
100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
@@
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
D.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$. We use a weighted CUSUM detector with RIPr defined in (4.17) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 6 to construct confidence sets. We report the average size and the conditional and unconditional coverage across $100$ independent runs in Table D.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
Table D.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 6 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
@@ D.3.2 Composite pre- and post-change (Setting III)
100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.94 0.94 23.32
@@
500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 23.04
Appendix E Another heuristic method for composite pre- or post-change
@@
For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section D.3.1 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1. Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section D.3.2 and we report the average size and the conditional and unconditional coverage across $100$ independent runs in Table E.1. From Table E.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
Table E.1: Setting II: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (3.15) with threshold $A=1000$.
@@
100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.94 0.94 25.83
@@
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.90 27.17
Table E.2: Setting III: Average (of $100$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (4.17) with threshold $A=1000$.
@@
100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 29.64
@@
500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$$\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 30.49
While we do not have any theoretical guarantees for this method, it is computationally more efficient than the previous approaches with theoretical guarantees, as demonstrated in Table E.3.
Summary of comparisons of heuristic and non-heuristic methods
We first compare the execution times of the three types of algorithms and present the results in Table E.3 for the Gaussian mean shift setting. Algorithm 3 and Algorithm 4 are collectively referred to as Proposed Method, which comes with theoretical guarantees but requires more computational time compared to the heuristic approaches. The first heuristic approach, described in Algorithm 5 and Algorithm 6, is referred to as Heuristic 1. Finally, the heuristic introduced in this section (which is Algorithm 2, with $F_{0}$ and $F_{1}$ being the “closest” pair of distributions in the composite pre- and post-change models) is labeled as Heuristic 2.
Table E.3: Average (of $10$ independent runs) execution times (in seconds) taken by different methods, with $N=B=100$, $\alpha=0.05$, and for the proposed method, $\beta=\gamma=0.05$, having pre- and post-change data-generating distributions as $N(0,1)$ and $N(1,1)$ respectively.
T A
Pre-change model
Post-change model
Proposed Method Heuristic 1 Heuristic 2
100 1000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 41.25 10.43 9.72
100 10000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 43.14 11.38 10.97
100 1000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 66.13 11.49 11.76
100 10000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 69.47 11.86 11.52
500 1000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 364.18 143.98 148.08
500 10000 $N(0,1)$$\{N(\mu,1):\mu>0.9\}$ 380.46 155.89 144.75
500 1000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 535.82 171.13 167.46
500 10000 $\{N(\mu,1):\mu\leq 0.1\}$$\{N(\mu,1):\mu>0.9\}$ 561.74 177.28 169.41
From Table E.3, we observe that the computational times are lower for the heuristic algorithms. Moreover, comparing the corresponding tables, we observe that the theoretically valid method (Proposed Method) gives higher coverage, also higher size of the confidence set, as compared to both heuristic methods. This suggests that while the heuristic approaches offer practical benefits in terms of speed, they may lead to less conservative inference, despite not having theoretical guarantees. This points at a natural direction for future work: to establish guarantees for our heuristic algorithms, or demonstrate settings where they empirically do not have coverage.

### v3

@@ H.3.1 Known pre-change and composite post-change (Setting II)
[…]
@@
[…]
H.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500. We use a weighted CUSUM detector with RIPr defined in (C.12) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. We report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table H.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
Table H.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
@@ H.3.2 Composite pre- and post-change (Setting III)
[…]
@@
[…]
Appendix I Another heuristic method for composite pre- or post-change
@@
For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section H.3.1 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1. Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section H.3.2 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1. From Table I.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
Table I.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.
@@
[…]
@@
[…]
Table I.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
@@
[…]
@@
[…]